If Soybean Meal is 16% protein and cornmeal is 8% protein how many pounds of each would you use to get 320 pounds of mixture that is 15% protein?
Answer by John10(297)

Let s be the amount of Soybean and c is the amount of cornmeal
The total is 320 lbs: s + c = 320
The mixture is 0.16s + 0.08c = 320(0.15)= 48
Solve this system:
s + c = 320
0.16s + 0.08c = 48
Hint: s = 280 lbs and c = 40 lbs
There you go!
If you need help with math 116/117 (algebra 2),let me know! John10